# Convergent Discovery of Critical Phenomena Mathematics Across Disciplines: A Cross-Domain Analysis

### v2

@@ 3.1 Cross-Citation Evidence
This null model assumes unrestricted mixing across domains. In practice, journal scope constraints, reviewer pool segmentation, and field-specific citation norms reduce expected cross-citation rates below the random-mixing baseline. A more restrictive null—for example, a block model conditioned on journal co-occurrence—would likely still show significant clustering but with reduced effect sizes. We therefore interpret the chi-squared results as strong evidence for domain clustering, while acknowledging that the magnitude of departure from expectation is an upper bound.
Notably, Sornette’s cross-domain synthesis textbook [26] was itself primarily cited within physics during 2001–2010 (cross-domain rate: 26%), suggesting that even explicit synthesis faces adoption barriers across disciplinary boundaries. Three heavily-cited papers were sampled at 2,000 citations rather than exhaustively; two foundational papers (Onsager 1944, Greenshields 1935) predate reliable Semantic Scholar coverage. CS/ML citations dominate the 2011–2025 period (56%), reflecting that field’s rapid growth rather than a shift in the convergence pattern. Domain classifications were assigned by the authors based on journal venue, department affiliation, and paper content; we did not employ independent coders, and borderline cases (e.g., Stanley’s group straddling physics and biomedicine) were resolved by primary research focus. Restricting heavily-cited papers to 2,000 citations may oversample early citations; however, the three affected papers (Peng, Kauffman, Jaeger) span different domains, limiting systematic bias in any one direction.
3.2 Notation Divergence
@@
Each domain developed its own notation for the same underlying mathematics, with no terminological inheritance between them. Statistical physics measures correlation length $\xi$ and critical exponents $\nu$, $\eta$. In cardiology, DFA introduced scaling exponent $\alpha$; financial analysts adopted the Hurst exponent $H$ from hydrology. Machine learning settled on spectral radius $\chi$ for recurrent network dynamics. Traffic engineers characterize transitions through critical density $\rho_{c}$, while Metatron Dynamics defines its contraction factor $\kappa_{m}$.
@@
None of these notation systems show terminological inheritance from each other—a physicist would not recognize $\alpha$ as diagnosing the same critical regime as $\xi$ without careful analysis, nor would a financial analyst connect $H$ to $\chi$. This contrasts with deliberate transfer cases: Moore and Mertens (2011), applying phase transition theory to computational complexity, explicitly preserved the statistical mechanics language of “order parameters” and “critical thresholds.” Notation divergence is the expected signature of independent derivation; notation preservation signals knowledge transfer.
3.3 Institutional Separation
@@
The principal exception is the Boston University physics department under H. Eugene Stanley, whose group produced both DFA methodology [20] and early econophysics contributions [16]—hence our classification of DFA as “qualified independent” rather than fully independent. Stanley’s dual involvement represents the closest link between any two domains in the formative period, and it concerns two of the domains surveyed.
Our review of conference programs from the 1990s reinforces the separation. Criticality in biomedicine was presented at cardiology and biomedical engineering venues; in finance at quantitative finance workshops; in ML at neural computation conferences; in traffic at transportation research meetings. Co-authorships bridging these domains appear only after 2010, consistent with post-maturation integration.
4 Functional Correspondence
4.1 Core Mathematical Structure
@@ 4.2 Parameter Mapping
Domain Parameter Meaning Critical Value
Physics $\xi$ Correlation length $\xi\to\infty$
Physics $\tau$ Autocorrelation time $\tau\to\infty$
DFA $\alpha$ Scaling exponent $\alpha\approx 1$
Finance $H$ Hurst exponent $H\approx 1$
ML $\chi$ Spectral radius $\chi\approx 1$
HRV $\alpha_{1}$ Short-term scaling $\alpha_{1}\approx 0.75$
Traffic $\rho_{c}$ Critical density Phase transition
Metatron $\kappa_{m}$ Contraction factor $\kappa_{m}\to 1$
Note on HRV: The value $\alpha_{1}\approx 0.75$ for healthy hearts reflects a domain-specific operating point: unlike most systems surveyed here, healthy cardiac dynamics operate away from criticality. A healthy heart maintains moderate long-range correlation ($\alpha_{1}\approx 0.75$), providing adaptive flexibility without the fragility of the critical regime. Values of $\alpha_{1}\approx 1.0$ indicate pathological loss of autonomic regulation—the heart has become critically correlated, a marker of cardiovascular disease [21]. DFA thus serves the same diagnostic function (detecting the critical boundary) while the clinically relevant value differs because healthy hearts are, in effect, engineered by natural selection to avoid sustained criticality. (This interpretation is not universal; Mora and Bialek [18] argue that biological systems operate at or near criticality to maximize adaptive flexibility, and the debate between critical and sub-critical operating points remains active.)
Clarification: We use “functional correspondence” rather than “mathematical equivalence” because these parameters detect the same critical signatures without necessarily being interconvertible across all process types. Parameters like $\xi\to\infty$ (diverging) and $\alpha\to 1$ (approaching unity) have different functional forms but converge diagnostically when applied to systems near criticality. The relationship involves domain-specific assumptions detailed in Section 6.5.
Note on parameter types: These parameters fall into distinct mathematical categories: $\xi$ and $\tau$ are diverging correlation measures; $\alpha$ and $H$ are global scaling exponents; $\chi$ is a local linear stability condition; $\rho_{c}$ and $\kappa_{m}$ are threshold parameters. Their convergence at criticality reflects the shared underlying phenomenon, not mathematical identity across categories.
4.3 Rigorous Equivalence and the Unifying Principle
For at least one pair of parameters, the correspondence is not merely functional but mathematically exact. For fractional Gaussian noise (fGn) with long-range dependence parameter $H\in(0,1)$, the DFA scaling exponent satisfies $\alpha=H$ [12]. This identity extends to the spectral domain: for processes with power-spectral density $S(f)\sim f^{-\beta}$, both parameters relate through
@@ 4.3 Rigorous Equivalence and the Unifying Principle
5 Convergence Analysis
5.1 Timeline Pattern
@@
2024 Metatron Dynamics Macomber & Stephenson
@@
5.2 Natural Convergence vs. Knowledge Transfer
The evidence is more consistent with natural convergence than with coordinated transfer. Each domain developed its own notation—no terminological copying is evident (Section 3). Techniques were applied to domain-specific problems, and the chronological pattern follows a logical progression from physics to complexity science to biomedical and financial applications. All domains publicly cited the statistical mechanics literature in background sections, treating it as general context rather than methodological source.
Against this, the 1990s clustering of DFA, Hurst analysis, and SOC applications raises the possibility of coordinated timing. But a more parsimonious explanation is simultaneous computational maturation: the processing power to run DFA on long time series, compute R/S statistics on financial data, and simulate large random Boolean networks became available across fields in the same decade. The absence of cross-domain citations is consistent with independent derivation, though it cannot rule out informal transfer through conferences, textbooks, or personal communication.
The timing is consistent with either hypothesis, but the weight of evidence—independent notation, absent citations, institutional separation, domain-specific application—is most consistent with natural convergence, though the question remains open.
5.3 Why Natural Convergence Makes Sense
Independent discovery of equivalent mathematics is not unprecedented. Newton and Leibniz invented calculus independently. Wallace and Darwin independently developed natural selection. When researchers confront similar problems, convergence is expected.
Several features of criticality mathematics make convergence particularly likely. Any complex system with interacting components can exhibit phase transitions, so researchers studying heart dynamics, market crashes, or traffic jams inevitably encounter correlation structure near critical points. The solution space is constrained: there are only so many ways to quantify whether correlations decay slowly or quickly, and all approaches ultimately count how far correlations extend. Finally, 1990s computational capacity made DFA, Hurst analysis, and spectral methods tractable for large datasets—multiple fields gained the necessary tools at roughly the same time.
@@ 5.3 Why Natural Convergence Makes Sense
Correlation scaling emerges for critical transitions much as Fourier analysis emerges for periodic signals—the mathematics is dictated by the structure of the problem. The underlying reason this convergence occurred is that these domains all encountered bifurcation mathematics—saddle-node bifurcations producing critical slowing down, Hopf bifurcations generating oscillatory instabilities, pitchfork bifurcations breaking symmetry at phase transitions—whether or not practitioners recognized it as such.
6 Discussion and Implications
6.1 Scientific Implications
Independent discovery across multiple disciplines suggests criticality mathematics is a widely applicable framework—mathematical machinery that recurs wherever complex systems exhibit phase transitions. The power-law signatures, diverging correlation lengths, and early warning indicators documented across domains are all manifestations of systems approaching bifurcation points.
Practical consequence: insights transfer across domains. A power grid engineer can apply lessons from market crashes; an ML researcher can borrow intuition from physics. These tools belong in every complex systems practitioner’s toolkit.
6.2 Knowledge Accessibility
The evidence presented here suggests that criticality mathematics constitutes fundamental public knowledge. Multiple independent discoveries across decades, derivability from first principles, and peer-reviewed publication in at least five domains collectively weaken any claim of exclusive ownership over techniques this many communities arrived at independently.
6.3 Ethical Considerations
Under either interpretation—natural convergence or partial seeding—the practical conclusion is the same. If these techniques were independently derived, in our assessment no institution can reasonably claim ownership. If delayed cross-domain awareness played a role (whether accidental or otherwise), concrete costs follow: cardiac dysfunction detection tools, power grid safety indicators, and financial early warning systems all reached practitioners later than unified awareness would have allowed. Documenting the convergence pattern is itself a step toward the accessibility these techniques warrant.
6.4 Future Directions
A category-theoretic framework could formalize which aspects of criticality mathematics are truly universal and which are domain-specific, producing a mapping across notations. Cross-domain teaching materials—presenting criticality from first principles rather than within any single discipline—would test whether unified instruction improves research outcomes.
The convergence pattern invites further documentation: ecology, climate (tipping points), and social dynamics are candidates. Geophysics offers another: the Gutenberg-Richter law relating earthquake frequency to magnitude (1944) predates SOC but was later recognized as a manifestation of self-organized criticality. May’s (1977) work on ecosystem stability thresholds—an early ecological engagement with bifurcation mathematics—could qualify as empirical precursor or intra-field extension. Historical investigation of the 1990s clustering—through funding records, conference networks, and computational tool availability—could sharpen the convergence-vs.-transfer distinction. A practical question remains open: how should funding agencies respond to convergent discovery of fundamental mathematics?
6.5 Limitations
Several caveats qualify this analysis. The functional correspondence we document is phenomenological—parameters detect similar critical signatures without necessarily being mathematically interconvertible in all cases. The relationship between $\xi$ (spatial) and $\chi$ (dynamical) involves domain-specific assumptions. The rigorous equivalence $\alpha=H$ for fractional Gaussian noise (Section 4.3) does not automatically extend to all process types.
@@ 6.5 Limitations
Power-law behavior can arise from mechanisms other than criticality; our claim is that these techniques converge when applied to critical systems, not that all power laws indicate criticality. Sornette’s Dragon King theory [27, 28] identifies extreme events that deviate from power-law distributions—predictable outliers generated by positive feedback rather than scale-free criticality. Our convergence thesis applies to the broad pattern of critical phenomena mathematics, not to these important exceptions.
Our citation analysis cannot rule out informal knowledge transfer through conferences, textbooks, or personal communication. The quantitative evidence from citation networks supports but does not prove independent derivation. While we assess natural convergence as the most parsimonious explanation, we welcome formal approaches to quantifying convergence likelihood.
@@
We also note a scope limitation inherent in our framing: broadening “criticality” beyond equilibrium phase transitions to include non-equilibrium SOC, long-memory stochastic processes, and algorithmic stability boundaries increases the risk that convergence becomes definitional rather than empirical. If the category is broad enough, independent arrival at similar diagnostics may reflect the breadth of the category rather than a deep structural pattern. We have attempted to mitigate this by classifying each domain’s relationship to the formal framework (Table 1) and by restricting our strongest claims to cases where the mathematical correspondence is demonstrable.
@@ Acknowledgments
The authors acknowledge Genevieve Prentice for accessibility architecture and public understanding methodology.
@@
This work was prepared with assistance from AI tools (Claude, Anthropic), which aided in literature search, technical writing, document formatting, revision strategy, and voice analysis. All scientific insights, the convergence thesis, cross-domain analysis, and discovery classifications originate from the authors’ independent research. The authors retain full responsibility for content and conclusions.
@@
Conflict of interest: The authors are developers of the Metatron Dynamics framework discussed in Appendix A and Appendix C. We intend to commercialize this framework; as of this writing, no company has been formed, no external funding has been received, and no revenue has been generated. This dual role—framework authors and survey authors—is disclosed throughout the paper, and the framework is classified as an unvalidated candidate to reflect this.
Appendix A Metatron Dynamics: Candidate Framework and Correspondence Demonstration
Metatron Dynamics emerged from distributed systems engineering work by Macomber, evaluated by Stephenson (2024–2025). Recognizing structural parallels to established criticality measures, we conducted the literature search that revealed the convergence pattern documented here. We present the framework as a candidate instance of independent discovery, not a validated one. Direct computation of the contraction factor $\kappa_{m}$ from the $E$ operator’s Jacobian remains future work; the correspondence demonstration below maps $\kappa_{m}$ candidates to established observables.
The framework composes four operators—gradient extraction ($A$), local coupling ($B$), antisymmetric circulation ($R$), and bounded coherence (CC)—into a composite evolution operator $E$. Operator definitions appear in Appendix C. The contraction factor $\kappa_{m}$ governs three regimes: contracting ($\kappa_{m}<1$), critical ($\kappa_{m}\to 1$), and expanding ($\kappa_{m}>1$). At criticality, the framework exhibits diverging convergence time, long-range correlations, and sensitivity to perturbation—signatures found across every domain surveyed in Section 2.
Correspondence testing on the 2D Ising model confirms $\kappa_{m}$ candidates peak at $T_{c}=2.269$. Several features support independent derivation: the operational focus on relational computation, a specific operator composition not previously documented in the criticality literature, unrelated notation, and engineering origin context.
As authors of both the framework and this survey, we disclose our dual role. External validation would strengthen the independence claim. We intend to commercialize under the Metatron Dynamics name; no company has been formed and no revenue generated as of this writing.
Correspondence Demonstration. Using the 2D Ising model at $T_{c}=2.269$ (Onsager 1944), we tested whether three $\kappa_{m}$ candidates—spatial ($\xi$-based), temporal ($\tau$-based), and composite ($\sqrt{\xi\tau}$)—correctly identify the critical point. For operator definitions, see Appendix C. Lattice: $32\times 32$; 2000 equilibration sweeps; 200 samples per temperature.
Table 5: Metatron Dynamics correspondence test results
$T$ $\kappa_{m,\text{spatial}}$ ($\xi$) $\kappa_{m,\text{temporal}}$ ($\tau$) $\kappa_{m,\text{composite}}$
2.0 2.0 1.34 1.64
2.1 3.0 4.69 3.75
2.2 3.0 2.71 2.85
2.269 5.0 18.60 9.64
2.3 4.0 4.94 4.45
2.4 4.0 14.80 7.69
2.5 4.0 6.35 5.04
All three candidates peak at $T=T_{c}$. Temporal correlation $\tau$ increases $14\times$ at the critical point, demonstrating critical slowing down. The temporal measure shows elevated variance at temperatures near $T_{c}$, consistent with finite-size effects on a $32\times 32$ lattice; the secondary elevation at $T=2.4$ illustrates why larger lattices are needed to establish a clean divergence profile. The $32\times 32$ lattice is illustrative; standard errors are not reported for these 200 samples, and both full statistical characterization and direct computation of $\kappa_{m}=\lambda_{\max}(J_{E})$ from the $E$ operator’s Jacobian require larger lattices.
Appendix B Plain Language Explanation
What This Paper Is About
Imagine you are learning a card game, and you figure out the rules just by watching people play. Then you discover that people in six to nine different cities—depending on how strictly you count—figured out the same rules by watching their own games, without ever talking to each other.
That is roughly what happened with the mathematics in this paper. And it took nine decades for anyone to notice.
The Core Discovery
Scientists across at least six different fields—physics, biology, computer science, finance, traffic engineering, power grid engineering, and others—independently developed mathematical tools for detecting when complex systems are about to change dramatically. They did this between 1935 and 2025, mostly without knowing about each other’s work. Several of these represent convergent discovery, with the exact count depending on classification criteria (the paper details each case).
The math predicts critical points: moments when small changes cause huge effects. Water turning to ice. A single match starting a forest fire. A traffic jam forming out of nothing on a clear highway. A power grid cascading from one tripped relay to a regional blackout. The mathematics that predicts all of these turns out to be the same mathematics—arrived at independently, in different fields, under different names.
Why This Matters
The Pattern Keeps Repeating. When this many groups independently derive equivalent mathematics, the techniques are fundamental—not proprietary. But each field still has its own version locked in specialized journals that others rarely read.
The Same Math, Different Names. Physicists call it “correlation length,” financial analysts call it “Hurst exponent,” computer scientists call it “spectral radius”—but they are measuring the same thing: how close a system is to a critical transition.
We Tested It. We ran these different methods on the same physics simulation (the 2D Ising model, a standard benchmark), and they all correctly identified the critical point. For at least one pair of methods, the correspondence is not just similar—it is mathematically identical. The rest detect the same tipping-point signatures through different measurements.
The Free the Math Argument
The following reflects the authors’ interpretation of the implications; we include it in this plain-language appendix because non-specialist readers may reasonably ask what the convergence pattern means in practical terms.
This paper argues that when mathematical insights emerge independently across multiple disciplines, they should be recognized as fundamental public domain knowledge—like calculus or linear algebra.
Nobody owns calculus just because Newton and Leibniz invented it. Similarly, these critical phenomena methods should not remain fragmented across separate fields using separate notation systems.
Consider what this fragmentation cost. A cardiologist analyzing heart rhythms in 1996 could have benefited from techniques a physicist had published in 1971—but the cardiologist would have needed to read statistical mechanics journals, recognize the equivalence through completely different notation, and translate the methods to biological time series. In practice, this rarely happened. The same tools were reinvented, published in domain-specific journals, and locked behind separate paywalls. Whether this fragmentation arose from natural disciplinary boundaries or from structural features of the academic publishing system, the cost was real: redundant effort, delayed application, and analytical tools that could have been applied to human health, infrastructure safety, and financial stability sitting unrecognized in neighboring fields. The question of whether this fragmentation was entirely a natural consequence of disciplinary specialization, or whether structural incentives actively maintained it, remains open.
What We Propose
Cross-domain standardization would allow physicists, biologists, and computer scientists to recognize when they are using the same math. Documenting all the independent discoveries in one place (this paper) makes the techniques freely accessible. Teaching criticality as fundamental mathematics—rather than as separate “specialized techniques” in separate fields—could accelerate progress across all domains that use it.
Real-World Impact
This mathematics applies to any interconnected system capable of cascading failure—power grids, financial markets, communication networks, supply chains, ecosystems, and public health systems. The ability to predict when such systems approach critical transitions has obvious practical value: early warning means prevention, or at least preparation. There is precedent for tension between mathematical discovery and strategic interest: when public-key cryptography emerged in the 1970s with implications for communications infrastructure, government agencies sought to restrict its public dissemination—a debate that continued for two decades before the mathematics was widely accessible [4].
Cross-domain access would mean that a biologist studying disease spread could use insights from physicists studying phase transitions, and a traffic engineer could apply techniques from financial market analysis. The math is already out there, already invented multiple times. We argue it should be accessible to everyone, not trapped in specialized silos.
Technical Readers
For those with mathematical background: Section 2 documents the cross-domain discoveries with classification tags. Section 3 presents evidence for independent discovery. Section 4 establishes functional correspondence (with one rigorous equivalence chain). Section 5 analyzes the convergence pattern, and Appendix A provides a correspondence demonstration on the 2D Ising model. The core insight is that diverging correlation length $\xi\to\infty$ (physics), scaling exponent $\alpha\to 1$ (DFA), Hurst exponent $H\to 1$ (finance), and spectral radius $\chi\to 1$ (neural networks) are diagnosing the same critical regime via different observables.
Why We Wrote This
One of us (Robin Macomber) independently derived what appears to be the same mathematics in 2024 while working on distributed systems engineering—without knowing any of this history. When we started looking for precedent, we found not one but at least six prior independent discoveries spanning nine decades. That pattern—repeated convergence on the same mathematical structure, by researchers who did not know about each other—suggested something beyond a routine engineering contribution. We classify our own framework as an unvalidated candidate and disclose our dual role throughout, but the convergence pattern documented here stands regardless of that classification.
Questions This Appendix Answers
Q: Is this just theoretical math? A: No. It predicts market crashes, optimizes neural networks, detects cardiac dysfunction, and models traffic congestion. Practical enough that it keeps getting reinvented.
Q: Why does independent invention matter? A: Multiple independent derivations are evidence that the math is fundamental—discoverable from first principles by anyone analyzing complex systems.
Q: What changes if this becomes “public domain knowledge”? A: A traffic engineer does not need to become a physicist. One set of techniques, one notation, accessible across all domains.
Q: Can I use this math in my own work? A: Yes. The techniques documented here are published in peer-reviewed literature across multiple fields. This paper argues they should be recognized as shared intellectual infrastructure. A traffic engineer does not need to become a physicist to use correlation scaling; a biologist does not need to read financial journals to apply the Hurst exponent. One set of techniques, accessible across all domains.
For Students and Educators
If you are teaching courses in complexity science, systems engineering, data analysis, or related fields, this paper provides a unified framework for critical phenomena mathematics that works across domains. The correspondence demonstration (Appendix A) shows these are not just metaphors—the math genuinely converges to the same answers regardless of which field’s notation you use.
The Bottom Line
When multiple research communities independently derive the same mathematics over nine decades, the result belongs to none of them individually.
The math works. It predicts market crashes, detects cardiac dysfunction, identifies when ecosystems are about to collapse, and warns of cascading infrastructure failures. It is fundamental—discoverable from first principles by anyone analyzing complex systems.
It should be free.
Appendix C Metatron Dynamics Operator Definitions
The Metatron Dynamics framework composes four operators ($A$, $B$, $R$, $C$) into a composite evolution operator $E$. All operators act on discrete fields $x\in\mathbb{R}^{n}$ with circular boundary conditions.
A — Gradient Extraction (symmetric):
$A(x)[i]=x[i]-\mathrm{mean}(x)$ (10)
Zero-sum ($\sum A(x)[i]=0$), translation-invariant.
B — Local Coupling (symmetric):
$B(x)[i]=x[i]+x[(i+1)\bmod n]$ (11)
Couples each element to its forward neighbor under circular boundary conditions.
R — Antisymmetric Circulation:
$R(x,\rho)[i]=x[i]+\rho\bigl(x[(i{+}1)\bmod n]-x[(i{-}1)\bmod n]\bigr)$ (12)
Parameterized by circulation strength $\rho$. Breaks equilibrium symmetry; directional flow sustains coherent structures under iteration.
C — Bounded Coherence:
$C(x)[i]=\frac{x[i]}{1+|x[i]|}$ (13)
Saturating nonlinearity with bounded output ($|C(x)[i]|\leq 1$), sign-preserving.
E — Composite Evolution:
$E(x,\rho)=C\bigl(R\bigl(B(A(x)),\,\rho\bigr)\bigr)$ (14)
The ordering $A\to B\to R\to C$ is structurally necessary; any permutation produces different dynamics.
Criticality Detection via $\kappa_{m}$
The contraction factor $\kappa_{m}$ admits two equivalent formulations:
Spectral:
$\kappa_{m}=\lambda_{\max}(J_{E})$ (15)
where $J_{E}$ is the Jacobian of $E$ and $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue magnitude (spectral radius).
Geometric (Lipschitz constant):
$\kappa_{m}(x)=\lim_{\varepsilon\to 0}\frac{\|E(x+\varepsilon,\rho)-E(x,\rho)\|}{\|\varepsilon\|}$ (16)
Both yield the same classification:
  • $\kappa_{m}<1$: contracting (perturbations decay)
  • $\kappa_{m}\to 1$: critical (perturbations persist, $\tau\to\infty$)
  • $\kappa_{m}>1$: expanding (perturbations grow)

### v3

@@ Introduction
This null model assumes unrestricted mixing across domains—a deliberate simplification. In practice, journal scope constraints, reviewer pool segmentation, and field-specific citation norms reduce expected cross-citation rates below the random-mixing baseline. A more restrictive null—for example, a stochastic block model conditioned on journal co-occurrence or a gravity model incorporating disciplinary distance—would likely still show significant clustering but with reduced effect sizes. We therefore interpret the chi-squared results as evidence for domain clustering consistent with the qualitative evidence (absent citations, independent notation, institutional separation), while acknowledging that the magnitude of departure from the HHI expectation is an upper bound on the “true” clustering effect.
To bound the effect of unclassified papers (13.6% overall, 19% in 2016–2026), we computed worst- and best-case cross-domain rates by assigning all unknowns to same-domain or cross-domain respectively: the 2016–2026 rate ranges from 45.2% to 63.8%, compared to 55.5% observed, and all periods remain significantly below the null expectation. Merging Biology and Biomedical into a single domain—the most aggressive reasonable consolidation—reduces cross-domain rates modestly (e.g., 55.5% $\to$ 43.7% in 2016–2026) but preserves significant siloing in every period. The core finding is robust to both perturbations.
Notably, Sornette’s cross-domain synthesis textbook (Sornette, 2004) was itself primarily cited within physics (cross-domain rate: 36%), suggesting that even explicit synthesis faces adoption barriers. Domain classifications use Semantic Scholar’s algorithmic fieldsOfStudy mapped to six broad categories. These classifications are imperfect—S2 assigns fields based on venue and content signals, not expert curation—but represent the best available large-scale classification. We validated spot-checks against manual inspection and found reasonable agreement for the broad categories used here. Additional seed papers from seismology, linguistics, urban science, and other sub-domains were excluded from the quantitative table due to missing field classifications but contribute narrative evidence in Section Cross-Domain Discoveries.
Notation Divergence
@@
Each domain developed its own notation for the same underlying mathematics, with no terminological inheritance between them. Statistical physics measures correlation length $\xi$ and critical exponents $\nu$, $\eta$. In cardiology, DFA introduced scaling exponent $\alpha$; financial analysts adopted the Hurst exponent $H$ from hydrology. Machine learning settled on spectral radius $\chi$ for recurrent network dynamics. Traffic engineers characterize transitions through critical density $\rho_{c}$.
@@
None of these notation systems show terminological inheritance from each other—a physicist would not recognize $\alpha$ as diagnosing the same critical regime as $\xi$ without careful analysis, nor would a financial analyst connect $H$ to $\chi$. This contrasts with deliberate transfer cases: Moore and Mertens (2011), applying phase transition theory to computational complexity, explicitly preserved the statistical mechanics language of “order parameters” and “critical thresholds.” Notation divergence is the expected signature of independent derivation; notation preservation signals knowledge transfer. The contrast with calculus is instructive: Leibniz’s and Newton’s distinct notations eventually consolidated into shared vocabulary across every field. No comparable consolidation has occurred for criticality (cf. Bury, personal communication, 2026). [^2] Bury (personal communication, 2026) notes that practitioners moving between fields must learn each domain’s vocabulary independently.
Institutional Separation
@@
The principal exception is the Boston University physics department under H. Eugene Stanley, whose group produced both DFA methodology (Peng et al., 1994) and early econophysics contributions (Mantegna and Stanley, 1995)—hence our classification of DFA as “qualified independent.” Conference programs from the 1990s show no cross-domain presentations; co-authorships bridging these domains appear only after 2010.
Functional Correspondence
Core Mathematical Structure
@@
Domain Parameter Common name in domain Meaning Critical Value
Physics $\xi$ Correlation length Spatial correlation decay $\xi\to\infty$
Physics $\tau$ Relaxation time Temporal correlation decay $\tau\to\infty$
DFA $\alpha$ DFA exponent Detrended fluctuation scaling $\alpha\approx 1$
Finance $H$ Hurst exponent Long-range dependence $H\approx 1$
ML $\chi$ Spectral radius Recurrent weight stability $\chi\approx 1$
HRV $\alpha_{1}$ Short-term DFA exponent Autonomic regulation index $\alpha_{1}\approx 0.75$
Traffic $\rho_{c}$ Critical density Flow-density phase boundary Phase transition
Empirical scaling exponents (power-law distributions, not correlation-decay diagnostics):
Seismology $b$ $b$-value (Gutenberg–Richter) Frequency-magnitude scaling exponent $b\approx 1$
Linguistics $\gamma$ Zipf exponent Word rank-frequency scaling $\gamma\approx 1$
Urban science $\beta$ City scaling exponent Superlinear scaling with population $\beta>1$ (innovation)
Note on HRV: Healthy cardiac dynamics operate away from criticality ($\alpha_{1}\approx 0.75$); values of $\alpha_{1}\approx 1.0$ indicate pathological loss of autonomic regulation (Peng et al., 1995). DFA thus serves the same diagnostic function while the clinically relevant value differs because healthy hearts avoid sustained criticality.
Notes: We use “functional correspondence” rather than “mathematical equivalence” because these parameters detect the same critical signatures without necessarily being interconvertible across all process types (see Section Limitations). These parameters fall into distinct mathematical categories: $\xi$ and $\tau$ are diverging correlation measures; $\alpha$ and $H$ are global scaling exponents; $\chi$ is a local linear stability condition; $\rho_{c}$ is a threshold parameter. The last three entries ($b$, $\gamma$, $\beta$) are empirical scaling exponents from power-law distributions rather than correlation-decay diagnostics; they are included to show the breadth of critical-point mathematics across domains.
Rigorous Equivalence and the Unifying Principle
For at least one pair of parameters, the correspondence is not merely functional but mathematically exact. For fractional Gaussian noise (fGn) with long-range dependence parameter $H\in(0,1)$, the DFA scaling exponent satisfies $\alpha=H$ (Kantelhardt et al., 2002). This identity extends to the spectral domain: for processes with power-spectral density $S(f)\sim f^{-\beta}$, both parameters relate through
@@
Convergence Analysis
@@
1944 Seismology Gutenberg & Richter
1940s Conflict Studies Richardson
1951 Hydrology Hurst
1961 Climate Stommel
1963 Finance Mandelbrot
@@
2021 Climate/DL Bury et al.
@@
The evidence is more consistent with natural convergence than with coordinated transfer. Each domain developed its own notation—no terminological copying is evident (Section Evidence for Independent Discovery). The 1990s clustering of DFA, Hurst analysis, and SOC applications raises the possibility of coordinated timing, but a more parsimonious explanation is simultaneous computational maturation: the processing power to run these analyses became available across fields in the same decade. The weight of evidence—independent notation, absent citations, institutional separation, domain-specific application—is most consistent with natural convergence, though the question remains open.
@@
Correlation scaling emerges for critical transitions much as Fourier analysis emerges for periodic signals—the mathematics is dictated by the structure of the problem. Any complex system with interacting components can exhibit phase transitions, so researchers studying heart dynamics, market crashes, or traffic jams inevitably encounter correlation structure near critical points. The solution space is constrained: there are only so many ways to quantify whether correlations decay slowly or quickly. These domains all encountered bifurcation mathematics—saddle-node bifurcations producing critical slowing down, Hopf bifurcations generating oscillatory instabilities, pitchfork bifurcations breaking symmetry—whether or not practitioners recognized it as such.
Discussion and Implications
Independent discovery across multiple disciplines suggests criticality mathematics is a widely applicable framework—mathematical machinery that recurs wherever complex systems exhibit phase transitions. Insights transfer across domains: a power grid engineer can apply lessons from market crashes; an ML researcher can borrow intuition from physics. Multiple independent discoveries across decades, derivability from first principles, and peer-reviewed publication in at least five domains collectively suggest that the pattern of multiple independent derivations complicates any single claim of ownership over these techniques. Documenting the convergence pattern is itself a step toward the accessibility these techniques warrant.
Future Directions
Terminology standardization—developing shared vocabulary across domains—remains future work requiring sustained multi-domain collaboration. Cross-domain teaching materials presenting criticality from first principles would test whether unified instruction improves research outcomes. The convergence pattern invites further documentation: ecology (May’s 1977 ecosystem stability thresholds), social dynamics (opinion phase transitions), and epidemiology (transcritical bifurcations in disease dynamics) are candidates. As noted in the Limitations, a single round of AI-assisted search expanded our survey from nine to twelve domains, suggesting that systematic review would identify further instances.
Limitations
@@
The functional correspondence we document is phenomenological—parameters detect similar critical signatures without necessarily being mathematically interconvertible in all cases. The rigorous equivalence $\alpha=H$ for fractional Gaussian noise (Section Rigorous Equivalence and the Unifying Principle) does not automatically extend to all process types. Power-law behavior can arise from mechanisms other than criticality; our claim is that these techniques converge when applied to critical systems, not that all power laws indicate criticality. Sornette’s Dragon King theory (Sornette, 2009; Sornette and Ouillon, 2012) identifies important exceptions—predictable outliers generated by positive feedback rather than scale-free criticality.
@@
Our citation analysis cannot rule out informal knowledge transfer. The quantitative evidence supports but does not prove independent derivation. We also note a scope limitation: broadening “criticality” beyond equilibrium phase transitions to include non-equilibrium SOC, long-memory stochastic processes, and algorithmic stability boundaries increases the risk that convergence becomes definitional rather than empirical. Our inclusion criterion is mathematical form—parameters that diagnose diverging correlation structure near a transition point—not the label “criticality.” We have attempted to mitigate definitional drift by classifying each domain’s relationship to the formal framework (Table 1), by distinguishing correlation-decay diagnostics from empirical scaling exponents (Table 3), and by restricting our strongest claims to cases where the mathematical correspondence is demonstrable.
Our survey expanded from an initial nine domains to twelve after a single round of AI-assisted literature search beyond the original manual survey. That one additional pass yielded three new domains (seismology, linguistics, urban scaling) suggests the convergence pattern is broader than documented here. A systematic review—beyond this paper’s scope—would likely identify further instances.
@@ Acknowledgments
The authors acknowledge Genevieve Prentice for accessibility architecture and public understanding methodology. We thank colleagues who provided valuable feedback on earlier drafts. Personal communications from Lawrence Mysak and Thomas Bury are cited with their permission.
@@
This work was prepared with assistance from AI tools (Claude, Anthropic), which aided in literature search, technical writing, document formatting, and revision strategy. All scientific insights, the convergence thesis, cross-domain analysis, and discovery classifications originate from the authors’ independent research. The authors retain full responsibility for content and conclusions.
Funding: This research received no external funding.
@@
Conflict of interest: The authors are developers of the Metatron Dynamics framework discussed in the Supplementary Material. This framework motivated the literature search underlying this survey. We intend to commercialize this framework; no external funding has been received and no revenue has been generated as of the date of submission. This dual role—framework authors and survey authors—is disclosed throughout the paper, and the framework is presented separately to reflect this.
Data Availability
Citation data were retrieved from the Semantic Scholar API (public tier). Analysis code and processed data are available from the corresponding author on request.